OpenStudy (anonymous):

4(1-3x)+1=-7 Can anyone please help?

OpenStudy (anonymous):

x=1 do u need an explanation?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

alright first you distribute the 4 to 1-3x, so that would be 4-12x+1=-7 then put like terms together. so 4+1= 5 and that would look like: 5-12x=-7 subtract 5 from both sides then it would be: -12x=-12, divide -12 by both sides and you end up with positive 1

OpenStudy (anonymous):

Thank you so much, you explained it perfectly.

OpenStudy (anonymous):

no problem :) message me if you need anymore help